OpenStudy (anonymous):

Solve the equation 25-4b=-6b+47 for b

OpenStudy (anonymous):

move -6b over to the other side and 25 to the other side, you will get, 2b = 22, then divide by 2, b = ?

OpenStudy (anonymous):

Isolate variable "b" to one side and your constants on the other. 25-4b=-6b+47 6b-4b=47-25 2b=22 Now when you divide both sides by 2 to isolate "b"... b=....